# Position paper

### v4

@@ 1 Introduction
Despite the impressive advances of generative AI (Cao et al., 2025), especially for natural language (large language models), vision (vision language models) and code (code models), recent investigations have pointed out a lack of competence in dealing with complex generation problems that require intricate planning (Kambhampati et al., 2024) and task adherence while keeping track of multiple constraints (Xie et al., 2024). A broad class of such complex problems, especially complex problems in expert domains, requires active human participation. Therefore, although the recent focus in generative AI has mostly been on complete automation (Hong et al., 2024; Brown et al., 2020), we believe that human-AI cooperation is a more promising approach.
To address complex problems of this kind, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction. In this article, we first give a general introduction to HAI-Co2 (Section 2), including its ethical challenges, and take first steps towards a formalization (Section 3). We outline open research challenges posed by HAI-Co2 in Section 4. Section 5 discusses related work. Section 6 concludes the paper.
@@ 2 HAI-Co2: Human-AI co-construction through preference-based search
Human-human cooperation would make less sense as a promising template for human-AI cooperation if current AI systems could solve complex expert-domain problems on their own. However, current AI capabilities are limited for complex expert-domain problems, e.g., due to insufficient knowledge and reasoning capabilities, bias and lack of trustworthiness. Scaling generative AI systems, particularly language models, has demonstrated improved performance across a range of tasks, e.g., math word problems and commonsense reasoning. However, even the most powerful LLMs show a lack of robustness under different semantic perturbations that would not have fooled an otherwise robust reasoner (Li et al., 2024). While it is unwise to rule out future improvements, their current limitations call for interventions beyond scaling. Thus, in order to be able to effectively solve complex expert-domain problems, we believe it is necessary for the AI agent to work closely with human experts.
@@
Is the current state of generative AI enough? As a running example, we will use a code generation example that illustrates some of the bottlenecks of GPT-4 Turbo, an update [^1] We use gpt-4-1106-preview. of the original GPT-4 model (OpenAI, 2023). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation is even more serious when more varied and realistic expressions of human preferences are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs struggle with handling complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgeable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unlock the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM, and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation.
@@ 3 Towards a formalization of HAI-Co2
As a running example, we provide the code generation use case in Figure 3 as an illustration of how different formal aspects of HAI-Co2 can be implemented (see Appendix A for more details). A second use case – in the expert domain of related work section generation – can be found in Appendix B.

### v5

@@ 1 Introduction
Despite the impressive advances of generative AI (Cao et al., 2025), especially for natural language (large language models), vision (vision language models) and code (code models), recent investigations have pointed out a lack of competence in dealing with complex generation problems that require intricate planning (Kambhampati et al., 2024) and task adherence while keeping track of multiple constraints (Xie et al., 2024a). A broad class of such complex problems, especially complex problems in expert domains, requires active human participation. Therefore, although the recent focus in generative AI has mostly been on complete automation (Hong et al., 2024; Brown et al., 2020), we believe that human-AI cooperation is a more promising approach.
To address complex problems of this kind, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction.
We acknowledge that many substrands of the NLP, ML, AI and HCI communities have noticed and addressed problems that are closely related to the problem HAI-Co2 is addressing. Moreover, many of the components of HAI-Co2 are also components of this prior work. The differentiator of our proposal lies (i) in the specifics of the problem statement, (ii) in the specifics of the holistic framework that we call HAI-Co2 and (iii) in a clear articulation of the research challenges and unmet needs that follow from the way we define problem and framework. We will lay out the research challenges in Section 4 and review related work in detail in Section 5. To set the context for the reader, we highlight here the most important points that make our work novel.
First, we address problems in expert domains that require complex solutions or artifacts.
  • We posit that an explicit and persistent representation of the solution space is crucial for systematic solution construction and that this solution space must be equipped to represent the complexities of expert domains, including through an abstraction hierarchy.
  • We propose search as the paradigm for constructing solutions: construction proceeds from an initial draft to a satisfactory solution step by step where each step consists of a search for an appropriate extension or modification of the current artifact.
Second, HAI-Co2 is designed for complex problem solving by a team of an expert and an AI agent.
  • In contrast to many other approaches, HAI-Co2 is set up for a cooperation of expert and agent as equal partners, each contributing their complementary strengths.
  • In complex domains, the exact goal of the cooperation is often underspecified in the beginning. This means that the cooperation is not only about constructing the solution, but also about constructing the precise objective of the solution. This raises ethical concerns (e.g., influencable reward functions) that need to be addressed.
  • We posit natural language as the primary medium of communication. One of the challenges in HAI-Co2 is that the agent has to learn effectively from implicit human feedback in natural language, but also from other signals in the complex co-construction environment, e.g., multimodal information and human edits.
  • Finally, a new evaluation methodology needs to be developed for HAI-Co2 due to the difficulty of assessing the quality of solutions in complex expert domains and due to the open-endedness and non-uniqueness of solutions to complex problems in expert domains.
In this article, we first give a general introduction to HAI-Co2 (Section 2), including its ethical challenges, and take first steps towards a formalization (Section 3). We outline open research challenges posed by HAI-Co2 in Section 4. Section 5 discusses related work. Section 6 concludes the paper.
@@ 2 HAI-Co2: Human-AI co-construction through preference-based search
Human-human cooperation would make less sense as a promising template for human-AI cooperation if current AI systems could solve complex expert-domain problems on their own. However, current AI capabilities are limited for complex expert-domain problems, e.g., due to insufficient knowledge and reasoning capabilities, bias and lack of trustworthiness. Scaling generative AI systems, particularly language models, has demonstrated improved performance across a range of tasks, e.g., math word problems and commonsense reasoning. However, even the most powerful LLMs show a lack of robustness under different semantic perturbations that would not have fooled an otherwise robust reasoner (Li et al., 2024b). While it is unwise to rule out future improvements, their current limitations call for interventions beyond scaling. Thus, in order to be able to effectively solve complex expert-domain problems, we believe it is necessary for the AI agent to work closely with human experts.
@@
Is the current state of generative AI enough? As a running example, we will use a code generation example that illustrates some of the bottlenecks of GPT-4 Turbo, an update [^1] We use gpt-4-1106-preview. of the original GPT-4 model (OpenAI, 2023). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024a)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation is even more serious when more varied and realistic expressions of human preferences are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs struggle with handling complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgeable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unlock the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM, and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation. On the other hand, forcing humans to take a passive role and making them review AI-generated code affects productivity negatively (Xu, Vasilescu, and Neubig, 2022; Bird et al., 2023). Meta-analyses by Simkute et al. (2024), with a focus on coding assistants, identify four major axes of AI-mediated productivity loss: shift of human roles from production to evaluation, unhelpful workflow restructuring, task interruptions, and, easy tasks becoming easier while hard tasks become harder. Recent advances in interactive coding have sought to address the first challenge by providing users with evaluation tools — AI-generated tests or static analysis-based — to decrease the cognitive load of reviewing AI-generated code. A key takeaway from these advancements is the need to redefine the human is supposed to do, what the AI needs to do, and how they are going to complement their respective expertise and limitations.
@@ 3 Towards a formalization of HAI-Co2
As a running example, we provide the code generation use case in Figure 3 as an illustration of how different formal aspects of HAI-Co2 can be implemented (see Appendix A for more details). A second use case – in the expert domain of related work section generation – can be found in Appendix B. The choice of code generation as a use case is motivated by recent advancements in automation in this domain and the subsequent development of knowledge regarding the limitations of these advancements. On the other hand, related work generation is a natural use case for the scientific community, irrespective of the individual researcher’s particular domain.
HAI-Co2, as we envision, is not limited to these two choices of use cases. For example, in case of travel planning problems (Xie et al., 2024b), one can conceptualize an abstract hierarchy similar to multi-agent planning systems (Li et al., 2025) and a plan-space search (Lee et al., 2025) guided by human preferences. Objective co-construction under HAI-Co2 allows the AI and the human to define the utility criteria of a plan together; e.g., an AI agent with access to large number of online reviews might inform the human about certain emergent patterns in hotel booking refusals, whereas an expert travel agent might guide the AI to book trips via certain transports. Similar analogies can be drawn for tasks pertaining to medical decision making. Recent studies (Tikhomirov et al., 2024) have indicated inherent differences in the clinical reasoning process adopted by humans and AI systems. This further strengthens our argument for the goal of AI to become complementary to human, not mimicking humans. One can formalize the solution space as space of concrete diagnosis, with the abstractions highlighting different aspects of the diagnosis that construct a hierarchical plan Fdez-Olivares et al. (2011).
@@ 4 Challenges and future research
Integration of vision-language models. The rapid advancement of vision-language models (VLMs) (Li et al., 2024a) presents opportunities for extending how human experts communicate preferences within HAI-Co2 beyond natural language. While visual representations in construction spaces (e.g., UML diagrams) are already achievable through text-based specifications and external rendering, VLMs could expand this ability and also enhance how experts communicate preferences about solutions at any abstraction level. VLMs would enable preference extraction from visual annotations, sketches, and spatial manipulations – providing an additional channel for the expert to convey utility information, particularly useful in inherently visual domains such as UI development. The key challenges mirror those already present in the text domain but with distinct complexities: (i) defining neighborhood structures for visual construction spaces requires formalizing continuous visual operations into discrete semantic transformations, and (ii) visual feedback signals must be decoded into the same utility learning framework that processes textual preferences, ensuring consistency across modalities. Future research should investigate how HAI-Co2’s mechanisms – the informational state, preference decoder, and hierarchical consistency constraints – can be extended to incorporate visual feedback while maintaining the systematic search methodology central to the framework.